# Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework

### v1

@@ 5.5 Are the Monetary Policy Shocks a Key Driver for Shaping the Macroeconomic Stars?
Given the meaningful effects of monetary policy shocks on the macroeconomic stars, an important question is whether these shocks have played a important role in shaping the historical dynamics of the stars. To address this, we conduct a counterfactual analysis based on the historical decomposition of the trend components, as shown in Figure 4.
@@
Figure 4: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
@@
Figure 4 compares the actual trend paths with counterfactual paths, which are generated by shutting down all identified monetary policy shocks, following the methodology outlined in Kilian and Lütkepohl, (2017). The shaded regions represent 68% credible bands around the counterfactual estimates.
@@
Our counterfactual analysis quantifies the contributions of monetary policy shocks over time. For instance, in Figure 4, Panel (a) shows that, absent these shocks, $\Delta g^{*}$ would have been over 1 percent higher during the crisis; Panel (b) indicates that $g^{*}$ would have been persistently higher, with a cumulative gap of about 0.5 percent over the sample period. Panel (c) shows that $\pi^{*}$, which remained below its counterfactual path since the early 1990s, would have hovered closer to 3 percent in the absence of repeated monetary tightening. Panel (d) shows that $r^{*}$ would have been more than 2 basis points higher during the financial crisis.
Taken together, the counterfactual analysis emphasizes that monetary policy is a main driver of the macroeconomic stars and tends to play a dual role: it effectively stabilizes long-run inflation expectations but can also impose a persistent drag on the real economy’s long-run trends when over-tightening occurs. Policymakers should therefore weigh the benefits of anchoring inflation against the potential costs for the real economy in the long run.
@@ 5.6 Robustness Analysis
We assess the robustness of our baseline findings by re-estimating the model under a range of alternative specifications. Figures C.1-C.8 show results of the IRFs of the trends and C.9-C.16 show the historical decomposition of all alternative specifications. Across all variations, the qualitative responses of the macroeconomic stars to a contractionary monetary policy shock remain similar to those in the baseline, indicating that our main results are not driven by specific modelling choices.
First, estimating the shrinkage parameter $V_{b}$ via a hierarchical Bayes approach yields posterior impulse responses that closely match the baseline, with slightly wider credible intervals (Figure C.1) and very similar historical decompositions (Figure C.9). Second, restricting the sample to pre-COVID observations ending in 2019Q4 produces similar negative responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ (Figure C.2) as well as an similar counterfactual path of the trends (Figure C.10), suggesting that the pandemic period does not drive the results. Third, increasing the VAR lag length from four to eight (Figure C.3 and igure C.11) does not materially alter the estimated responses or historical decompositions. Fourth, replacing the federal funds rate with the one-year or two-year Treasury yield (Figures C.4, C.5, C.12 and C.13) preserves the qualitative pattern of the responses and the pattern of the counterfactual pathes. Fifth, substituting the GDP deflator with the PCE price index (Figures C.6 and C.14) yields comparable estimates. Sixth, using the unadjusted monetary policy surprise measure from Bauer and Swanson (2023) instead of the orthogonalized version (Figures C.7 and C.15) results in nearly identical posterior distributions. Finally, controlling for financial conditions by adding the excess bond premium (Favara et al., , 2016) to the cycle part of the model confirms our main empirical findings( Figures C.8 and C.16).
Overall, these robustness analyses confirm that the estimated effects of monetary policy shocks on the macroeconomic stars are stable across alternative sample periods, lag specifications, interest rate measures, inflation measures, external instruments, and controlling for financial conditions using the excess bond premium.
@@ 6 Conclusions
This paper develops SMUC-IV to explore the effects of monetary policy shocks on the macroeconomic stars, measured as long-run trends, in a unified framework. Using our SMUC-IV, we show that monetary tightening can have negative effects on the growth rate of potential output, trend inflation, and the neutral interest rate, and that monetary policy shocks are an important driver of the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate. From a policy perspective, the results indicate that policymakers can use monetary policy to re-anchor long-run inflation expectations at target, but also caution that too much tightening could harm the real economy in the long run. Recently, growing attention has been paid to the long-run effects not only of monetary policy shocks but also of other structural shocks, such as financial, fiscal, and demand shocks (Cerra and Saxena, , 2005, 2008; Antolin-Diaz and Surico, , 2025; Furlanetto et al., , 2025). Against this backdrop, future research could extend our framework to examine how different types of structural shocks shape long-run macroeconomic equilibria.
@@
where $\bar{\bm{\eta}}_{t}=(\Delta g_{t}^{*}-\Delta g_{t-1}^{*},\pi_{t}^{*}-\pi_{t-1}^{*},r_{t}^{*}-r_{t-1}^{*},\mathbf{c}_{t}^{\prime},m_{t})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}^{\prime}$ and $\bar{\mathbf{A}}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i},0)$ for $i=1,\ldots,p$. Stacking equation (25) over $t=1,\ldots,T$, we have
@@
where $\bar{\bm{\eta}}=(\bar{\bm{\eta}}_{1}^{\prime},\ldots,\bar{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\bar{\mathbf{a}}=\text{vec}\left((\bar{\mathbf{A}}_{1},\ldots,\bar{\mathbf{A}}_{p})^{\prime}\right)$, $\bar{\mathbf{X}}=(\bar{\mathbf{X}}_{1}^{\prime},\ldots,\bar{\mathbf{X}}_{T}^{\prime})^{\prime}$ with $\bar{\mathbf{X}}_{t}=\mathbf{I}_{7}\otimes(\bar{\bm{\eta}}_{t-1}^{\prime},\ldots,\bar{\bm{\eta}}_{t-p}^{\prime})$. Let $\bm{\phi}=\text{vec}\left((\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})^{\prime}\right)$ and $\mathbf{S}_{\phi}$ be the selection matrix such that $\bar{\mathbf{a}}=\mathbf{S}_{\phi}\bm{\phi}$. Then we can write (26) as
@@
In this Appendix, we first present the analytical expressions for (21) and (22) and then detail the construction of the tuning function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$.
Expressions for $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ and $p(\bm{\Phi})$
From result in (20), let $\bm{\Lambda}=\mathbf{K}^{-1}$ and partition $\bm{\Lambda}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
The property of Gaussian distribution implies that the marginal distribution of $\mathbf{y}$ (of dimension $4T\times 1$) is a Gaussian with means $\bm{\mu}_{y}$ and covariance matrix $\bm{\Lambda}$. This yields
Next, we derive the expression for the marginal prior $p(\bm{\Phi})$. Given the priors of $\bm{\Phi}$ and $(\kappa_{1},\kappa_{2})$ described in Section 3.1, we have
@@
The last equality holds because the integrals in the third line are the kernels of inverse-gamma densities.
@@
Figure C.1: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the shrinkage parameter $V_{b}$ using a hierachical Bayes approach.
Figure C.2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we use only data unitl 2019Q4.
@@
Figure C.7: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
@@
Figure C.9: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the shrinkage parameter $\lambda_{B}$ using a hierachical Bayes approach.
@@
Figure C.10: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we use only data unitl 2019Q4.
@@
Figure C.11: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with eight lags instead of four.
@@
Figure C.12: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the one-year treasury yield instead of using the federal funds effective rate.
@@
Figure C.13: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the two-year treasury yield instead of using the federal funds effective rate.
@@
Figure C.14: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the personal consumption expenditure price index instead of the GDP deflator.
@@
Figure C.15: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
@@
Figure C.16: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with adding the excess bond premium.

### v2

@@ 5.5 The Role of Monetary Policy Shocks in Shaping Macroeconomic Stars
Given the meaningful effects of monetary policy shocks on the macroeconomic stars, an important question is whether these shocks have played an important role in shaping the historical dynamics of the stars. To address this, we conduct a counterfactual analysis based on the historical decomposition of the trend components, as shown in Figure 6.
@@
Figure 6: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
@@
Figure 6 compares the actual trend paths with counterfactual paths, which are generated by shutting down all identified monetary policy shocks, following the methodology outlined in kilian2017structural. The shaded regions represent 68% credible bands around the counterfactual estimates.
@@
Our counterfactual analysis quantifies the contributions of monetary policy shocks over time. For instance, in Figure 6, Panel (a) shows that, absent these shocks, $g^{*}$ would have been over 0.8 percentage points higher during the crisis. Panel (b) shows that $\pi^{*}$, which has remained below its counterfactual path since the early 1990s, would have hovered closer to 3.5 percent in the absence of monetary tightening. Panel (c) shows that $r^{*}$ would have been around 3 percentage points higher during the financial crisis.
5.6 Results Discussion and Policy Implications
The negative effects of contractionary monetary policy shocks on $g^{*}$ and $r^{*}$ shown in Figure 4 can be rationalized by the mechanism emphasized in the Keynesian endogenous growth model with nominal rigidities in fornaro2023scars. In that setting, long-run growth is endogenous and can be influenced by both demand- and supply-side forces. With nominal rigidities, output can deviate from potential, so monetary policy has real effects. In their framework, monetary tightening can trigger a “supply–demand doom loop”: a contraction reduces aggregate demand, compressing firms’ profits and weakening incentives to invest. The resulting decline in investment lowers expected productivity growth and household wealth, which further depresses demand and induces additional cutbacks in investment. As a result, the $g^{*}$ and $r^{*}$ fall. Moreover, by amplifying the effects of adverse disturbances on the real economy, this vicious spiral helps interpret panel (a) and (c) in Figure 6: while the actual paths of $g^{*}$ and $r^{*}$ (with contractionary monetary policy shocks) and the counterfactual paths (excluding the shocks’ effects) share a similar downward tendency, the actual series remain persistently below their counterfactual counterparts.
Turning to trend inflation, at first glance, significant negative effects of monetary policy shocks in panel (b) of Figure 4 may appear undesirable, as they could be read as evidence of de-anchoring in long-run inflation expectations. However, if long-run expectations are persistently off target, such effects may instead reflect a re-anchoring channel, whereby policy deliberately shifts expectations back toward the target (for detailed discussion, see, e.g., diegel2021long). Consistent with this interpretation, our counterfactual analysis (panel (b) of Figure 6) indicates that monetary policy shocks pulled trend inflation down and kept it closer to the 2% target from the mid-1990s through the onset of the COVID-19 pandemic.
Taken together, our results suggest that central banks face a dilemma: bringing inflation back to target may come at the cost of lower potential GDP growth and a lower neutral interest rate. This, in turn, points to a potential role for fiscal interventions that support business investment and the economy’s productive capacity during disinflation episodes.
@@ 5.7 Robustness Analysis
We assess the robustness of our baseline findings by re-estimating the model under a range of alternative specifications. Figures C.1-C.8 show results of the IRFs of the trends and Figures C.9-C.16 show the historical decomposition of all alternative specifications. Across all variations, the qualitative responses of the macroeconomic stars to a contractionary monetary policy shock remain similar to those in the baseline, indicating that our main results are not driven by specific modelling choices.
First, estimating the shrinkage parameter $V_{b}$ via a hierarchical Bayes approach yields posterior impulse responses that closely match the baseline, with slightly wider credible intervals (Figure C.1) and very similar historical decompositions (Figure C.9). Second, restricting the sample to pre-COVID observations ending in 2019Q4 produces similar negative responses of $g^{*}$, $\pi^{*}$, and $r^{*}$ (Figure C.2) as well as a similar counterfactual path of the trends (Figure C.10), suggesting that the pandemic period does not drive the results. Third, increasing the VAR lag length from four to eight (Figures C.3 and C.11) does not materially alter the estimated responses or historical decompositions. Fourth, replacing the federal funds rate with the one-year or two-year Treasury yield (Figures C.4, C.5, C.12 and C.13) preserves the qualitative pattern of the responses and the pattern of the counterfactual paths. Fifth, substituting the GDP deflator with the PCE price index (Figures C.6 and C.14) yields comparable estimates. Sixth, using the unadjusted monetary policy surprise measure from Bauer and Swanson (2023) instead of the orthogonalized version (Figures C.7 and C.15) results in nearly identical posterior distributions. Finally, controlling for financial conditions by adding the excess bond premium (favara2016updating) to the cycle part of the model confirms our main empirical findings( Figures C.8 and C.16).
Overall, these robustness analyses confirm that the estimated effects of monetary policy shocks on the macroeconomic stars are stable across alternative sample periods, lag specifications, nominal interest rate measures, inflation measures, external instruments, and controlling for financial conditions using the excess bond premium.
@@ 6 Conclusions
This paper develops SMUC-IV to explore the effects of monetary policy shocks on the macroeconomic stars, measured as long-run trends, in a unified framework. Using our SMUC-IV, we show that monetary tightening can have negative effects on the potential GDP growth, trend inflation, and the neutral interest rate, and that monetary policy is an important driver of these stars. From a policy perspective, our results indicate that policymakers can bring elevated inflation back to target through monetary tightening, but potentially at the cost of long-run damage to the real economy. This dilemma points to a role for complementary fiscal interventions that support business investment and the economy’s productive capacity during disinflation episodes. Recently, growing attention has been paid to the long-run effects not only of monetary policy shocks but also of other structural shocks, such as financial, fiscal, and demand shocks (cerra2005did; cerra2008growth; antolin2025long; furlanetto2025estimating). Against this backdrop, future research could extend our framework to examine how different types of structural shocks shape long-run macroeconomic equilibria.
@@
where $\bar{\bm{\eta}}_{t}=(g_{t}^{*}-g_{t-1}^{*},\pi_{t}^{*}-\pi_{t-1}^{*},r_{t}^{*}-r_{t-1}^{*},\mathbf{c}_{t}^{\prime},m_{t})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}^{\prime}$ and $\bar{\mathbf{A}}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i},0)$ for $i=1,\ldots,p$. Stacking equation (25) over $t=1,\ldots,T$, we have
@@
where $\bar{\bm{\eta}}=(\bar{\bm{\eta}}_{1}^{\prime},\ldots,\bar{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\bar{\mathbf{a}}=\text{vec}\left((\bar{\mathbf{A}}_{1},\ldots,\bar{\mathbf{A}}_{p})^{\prime}\right)$, $\bar{\mathbf{X}}=(\bar{\mathbf{X}}_{1}^{\prime},\ldots,\bar{\mathbf{X}}_{T}^{\prime})^{\prime}$ with $\bar{\mathbf{X}}_{t}=\mathbf{I}_{7}\otimes(\bar{\bm{\eta}}_{t-1}^{\prime},\ldots,\bar{\bm{\eta}}_{t-p}^{\prime})$. Let $\bm{\phi}=\text{vec}\left((\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})^{\prime}\right)$ and $\mathbf{S}_{\phi}$, then there is a selection matrix $\mathbf{S}_{\phi}$ such that $\bar{\mathbf{a}}=\mathbf{S}_{\phi}\bm{\phi}$. It follows that we can write (26) as
@@
In this Appendix, we first present the analytical expression of the marginal prior $p(\bm{\Phi})$ in (22), and then detail the construction of the tuning function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$.
Expressions for $p(\bm{\Phi})$
Given the priors of $\bm{\Phi}$ and $(\kappa_{1},\kappa_{2})$ described in Section 3.1, we have
@@
Figure C.1: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the shrinkage parameter $V_{b}$ using a hierarchical Bayes approach.
Figure C.2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we use only data until 2019Q4.
i
@@
Figure C.7: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (bauer2023reassessment) instead of using the orthogonalized monetary policy surprise measure.
@@
Figure C.9: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the shrinkage parameter $\lambda_{B}$ using a hierarchical Bayes approach.
@@
Figure C.10: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we use only data until 2019Q4.
@@
Figure C.11: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with eight lags instead of four.
@@
Figure C.12: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with the one-year treasury yield instead of using the federal funds effective rate.
@@
Figure C.13: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with the two-year treasury yield instead of using the federal funds effective rate.
@@
Figure C.14: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with the personal consumption expenditure price index instead of the GDP deflator.
@@
Figure C.15: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (bauer2023reassessment) instead of using the orthogonalized monetary policy surprise measure.
@@
Figure C.16: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural. In this plot we estimate the model with adding the excess bond premium.